Two years after an expansion, a company had sales of $42,000. Four years later the sales were $102,000. Assuming that the sales in dollars S and the time in years t are related by a linear equation, find the equation relating S and t.

Answer by stanbon(75887)

You have two points and want to find the equation of the
line through those points.
The points are (2,42000) and (6,102000)
The slope if 60000/4= 15000
Substitute into y=mx+b to find "b"
42000=(15000)(2)+b
12000=b
The equation you want is S=(15000)t+12000
Cheers,
Stan H.